Operations Report 23Nov2019


Crew 216 Operations Report 23-Nov-2019
SOL: -1
Name of person filing report: Shane Usher/Mike Lawson
Non-nominal systems: N/A
Notes on non-nominal systems: N/A
Generator: run
Hours run: 12 h -5 min
From what time last night: 1910
To what time this morning: 0705
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night): 72%
Diesel Reading – 45%
Station Propane Reading – 63%
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 47
Water Meter (units): 1460749
Water (static tank) (gal): 500
Static to Loft Pump used – yes
Water in Green Hab (gal): 225.7
Water in Science Dome (gal): 0
Toilet tank emptied: yes
Deimos rover used: no
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Sojourner rover used: Assigned to Director
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Spirit rover used: no
Hours: 112.7
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 100%
Currently charging: yes
Opportunity rover used: no
Hours: 69
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes
Curiosity rover used: no
Hours: 119.1
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 100%
Currently charging: yes
Notes on rovers: N/A
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used (gal): 0
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director
HabCar used and why, where? With Director
CrewCar used and why, where? At MDRS
General notes and comments: N/A
Summary of internet: Functional
Summary of suits and radios: N/A
Summary of Hab operations: N/A
Summary of GreenHab operations: Atmospheric equalization to the green hab carried out during the afternoon, plants were watered (16L,4.6 gal). Door closed at 5pm, heater on overnight.
Summary of Science Dome operations: N/A
Summary of RAM operations: N/A
Summary of any observatory issues: Nominal
Summary of health and safety issues: Broken smoke alarm in Greenhab will be replaced.
Questions, concerns and requests to Mission Support: N/A

Crew 216 Sol Summary 23Nov2019

[title Sol Summary – November 23rd]

[category sol-summary]

Crew 216 Sol Summary Report 23-NOV-2019

Sol: -1

Summary Title: Change over day

Author’s name: Marc Levesque

Mission Status: Inactive

Sol Activity Summary: Today was a hand over day with Crew 215 who introduced our crew to MDRS and its systems. We have learned engineering procedures, identified EVA locations, GreenHab operations, inquired about rover operations, but tomorrow is when all the familiarizations will be carried out.

Look Ahead Plan: Tomorrow we will be running through Science Dome operations, GreenHab operations, RAM operations, EVA procedures and equipment, rover operations, and familiarization of the Hab operations, including cooking, cleaning, and maintenance.

Anomalies in work: N/A.

Weather: Just freezing overnight, light winds, clear skies.

Crew Physical Status: Healthy


Reports to be file: Sol Summary Report, Operations Report

Support Requested: N/A

Copyright © The Mars Society. All rights reserved. | Main Site